a chemist wants to mix a 60% saline solution with 20 liters of a 45% saline solution to obtain a 55% saline solution. how much of the 60% saline solution should she use?
Answer by Earlsdon(6294)

Let x + the number of liters of the 60% saline solution.
The amount of salt in the solutions can be expressed by:
0.6x+0.45(20) = 0.55(x+20) Simplify.
0.6x+9 = 0.55x+11 Solve for x. Subtract 0.55x from both sides.
0.05x+9 = 11 Subtract 9 from both sides.
0.05x = 2 Divide both sides by 0.05
x = 40
The chemist should mix 40 liters of the 60% saline solution with 20 liters of 45% saline solution to obtain 60 liters of 55% saline solution.